How many 4-digit numbers can be formed from the Set A = (0,1,2,3,4,5,6) if there is no repetition.
My answer is n=6*6*5*4 = 720 4-DIGIT numbers can be formed.
How many of the numbers in part a are divisible by 5? (You must consider 2 cases.)
My answer is 720/5=144
what is the 2 cases I should be looking at?
Pls help

Answer by Edwin McCravy(20054)

How many 4-digit numbers can be formed from the Set A = (0,1,2,3,4,5,6) if there is no repetition.
My answer is n=6*6*5*4 = 720 4-DIGIT numbers can be formed.

That's right!

How many of the numbers in part a are divisible by 5? (You must consider 2 cases.)
My answer is 720/5=144

That's wrong!

what is the 2 cases I should be looking at?
Pls help

If a number is divisible by 5 it must end in either a 5 or a 0.

(0,1,2,3,4,5,6)

Choose the fourth digit first.

Case 1.  Choose the fourth digit as 0

That leaves 6 ways to choose the first digit.
That leaves 5 ways to choose the second digit.
That leaves 4 ways to choose the third digit. 

That's 6*5*4 = 120 ways

Case 2.  Choose the fourth digit as 5

That leaves 5 ways to choose the first digit. (It can't be 0)
That leaves 5 ways to choose the second digit.
That leaves 4 ways to choose the third digit. 

That's 5*5*4 = 100 ways 

That's a total of 120 + 100 or 220 ways a 4 digit 
number can end in 0 or 5.

Edwin